OpenStudy (anonymous):

Consider the line 2x-3y=-2 . What is the slope of a line parallel to this line? What is the slope of a line perpendicular to this line?

OpenStudy (kymber):

Parallel lines have the same slope and perpendicular lines have opposite reciprocal slopes. First you need to put it into y=mx+b form.

OpenStudy (anonymous):

y=2/3x+2/3

OpenStudy (anonymous):

Parallel lines will have the same slope, so put 2x-3y=-2 in y=mx+b form. So, solve for y in your equation. The m stands for slope.

OpenStudy (anonymous):

then what

OpenStudy (kymber):

M is the slope in y=mx+b. So follow what I just said about parallel and perpendicular lines to get the slopes.

OpenStudy (anonymous):

i typed in perpendicular as 3/2. is that right?

OpenStudy (anonymous):

it says its wrong

OpenStudy (kymber):

No, perpendicular is opposite /and/ the reciprocal! Reciprocal being what you did, and opposite meaning positive/negative. So it would be -3/2. :P